OpenStudy (anonymous):

what is 2x2-x+2xy- y?

OpenStudy (anonymous):

factor?

OpenStudy (anonymous):

Again, we need more information. DO you want us to factor it?

OpenStudy (anonymous):

\[2x ^{2}-x+2xy-y\] = x(2x-1)+y(2x-1) =(2x-1)(x+y)